Reports 1-1 of 1 Clear search Modify search
On-site interventions (General)
derossi, gosselin, bulten, chiummo, spinicelli - 17:24 Tuesday 06 April 2021 (51332) Print this report
Comment to INJ on safety mode and check DET ERoom Temperature (51314)

The system has been recovered this afternoon.

One of the actuators of the EIB suspensions was out of range and we called Henk Jan to recover it.

Then we had to manually realing the beam with EIB M4 and M8 mirrors in order to be able to close the beam pointing control loop. Once the BPC was closed, we set the old values for the SIB1 and MC positions but the IMC was not locked yet. The references for the MC had slighlty changed and we had to move a little bit MC TX in order to lock the mode cleaner (attached plot).

Images attached to this comment
Search Help
×

Warning

×